# Semi-analytical covariance matrices for two-point correlation function for DESI 2024 data

### v2

@@ A.1 Full
Analogously, and are the 3-point and connected 4-point correlation functions of the tracers listed in the superscript evaluated at the separations between and points, respectively. These non-Gaussian higher-point functions are included for completeness but nulled in practice, we reflected this by crossing them out in the expressions.
A.2 Jackknife
The full multi-tracer jackknife covariance model can be constructed analogously to Eq. A.1 but with the following terms:
where is an additional jackknife weight tensor:
@@ B.1 Angular bins
In pycorr, during the combination of regions labeled by , the total counts (where each can be or , data or randoms) become
Prior to that, all the counts are brought to the same normalization within each region, which does not affect the correlation function. The norms for the total counts of each type are sums of the corresponding norms for the two regions. So they are also the same for all total counts. Therefore, in all these cases, the correlation function is equal to the ratio of non-normalized counts, not only the normalized ones. Then the total correlation function is simply the average weighted by counts:
@@
where we neglect the covariance between the correlation functions in different regions, the estimation of which poses extra challenges. For the North and South Galactic caps in the DESI footprint, which are well separated from each other, this seems like a safe approximation.
@@ B.2 Legendre
One can do the reverse approximately with bin-averaged values of the Legendre polynomials171717Or possibly with just the middles of the bins, but this way they end up very much related to already computed projection factors .:
These do not depend on the region, and work equally well for the combined one as for parts. Additional approximation comes from the fact that we limit the multipole index — in this work we only consider .
So we can work out the partial derivative of the Legendre moment of the combined correlation function with respect to one in each of the regions in the following steps:
@@
Appendix C Shot-noise rescaling based on mocks
We show the procedure for fitting the shot-noise rescaling to the mock sample covariance schematically in Fig. 4. It is similar to the method of [26], used before utilizing jackknife was proposed in [27]. The resulting RascalC covariance matrix is the best fit (in terms of KL divergence) of the full covariance model to the mock sample covariance matrix; this procedure can be seen as theoretically motivated smoothing.
NGC | SGC | |||
Jackknife | Mock sample | Jackknife | Mock sample | |
LRG pre | ||||
LRG post | ||||
ELG pre | ||||
ELG post | ||||
BGS pre | ||||
BGS post |
We show the shot-noise rescaling values obtained from jackknife (the main method) and mock sample covariance in Table 10. The numbers are very close in all cases. The deviations from the mean are not highly correlated mock-to-mock, but this can be expected due to additional scatter in individual jackknife covariances – the mock sample covariance used for reference was based on all 1000 mocks and thus the same for 10 single-mock runs for each case.
Appendix D Author affiliations
Center for Astrophysics Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA
@@ Appendix D Author affiliations
Institute of Physics, Laboratory of Astrophysics, École Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, CH-1290 Versoix, Switzerland
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
Physics Department, Yale University, P.O. Box 208120, New Haven, CT 06511, USA
Department of Physics & Astronomy, Ohio University, Athens, OH 45701, USA
@@
Center for Cosmology and AstroParticle Physics, The Ohio State University, 191 West Woodruff Avenue, Columbus, OH 43210, USA
@@
Department of Astronomy, The Ohio State University, 4055 McPherson Laboratory, 140 W 18th Avenue, Columbus, OH 43210, USA
Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA
Physics Dept., Boston University, 590 Commonwealth Avenue, Boston, MA 02215, USA
University of Michigan, Ann Arbor, MI 48109, USA
@@
Leinweber Center for Theoretical Physics, University of Michigan, 450 Church Street, Ann Arbor, Michigan 48109-1040, USA
@@
Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK
@@
Institute for Computational Cosmology, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK
@@
Instituto de Física, Universidad Nacional Autónoma de México, Cd. de México C.P. 04510, México
@@
Department of Astronomy, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
@@
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Menlo Park, CA 94305, USA
SLAC National Accelerator Laboratory, Menlo Park, CA 94305, USA
University of California, Berkeley, 110 Sproul Hall #5800 Berkeley, CA 94720, USA
@@
Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra Barcelona, Spain
Departamento de Física, Universidad de los Andes, Cra. 1 No. 18A-10, Edificio Ip, CP 111711, Bogotá, Colombia
@@
Observatorio Astronómico, Universidad de los Andes, Cra. 1 No. 18A-10, Edificio H, CP 111711 Bogotá, Colombia
Department of Physics, The University of Texas at Dallas, Richardson, TX 75080, USA
@@
Departament de Física Quàntica i Astrofísica, Universitat de Barcelona, Martí i Franquès 1, E08028 Barcelona, Spain
Institut d’Estudis Espacials de Catalunya (IEEC), 08034 Barcelona, Spain
@@
Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (UB), c. Martí i Franquès, 1, 08028 Barcelona, Spain.
@@
Consejo Nacional de Ciencia y Tecnología, Av. Insurgentes Sur 1582. Colonia Crédito Constructor, Del. Benito Juárez C.P. 03940, México D.F. México
@@
Departamento de Física, Universidad de Guanajuato - DCI, C.P. 37150, Leon, Guanajuato, México
Fermi National Accelerator Laboratory, PO Box 500, Batavia, IL 60510, USA
@@
Department of Physics, The Ohio State University, 191 West Woodruff Avenue, Columbus, OH 43210, USA
School of Mathematics and Physics, University of Queensland, 4072, Australia
NSF NOIRLab, 950 N. Cherry Ave., Tucson, AZ 85719, USA
@@
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), FR-75005 Paris, France
@@
Departament de Física, Serra Húnter, Universitat Autònoma de Barcelona, 08193 Bellaterra (Barcelona), Spain
@@
Laboratoire de Physique Subatomique et de Cosmologie, 53 Avenue des Martyrs, 38000 Grenoble, France
@@
Institució Catalana de Recerca i Estudis Avançats, Passeig de Lluís Companys, 23, 08010 Barcelona, Spain
Department of Physics and Astronomy, University of Sussex, Brighton BN1 9QH, U.K
Department of Physics & Astronomy, University of Wyoming, 1000 E. University, Dept. 3905, Laramie, WY 82071, USA
@@
National Astronomical Observatories, Chinese Academy of Sciences, A20 Datun Rd., Chaoyang District, Beijing, 100012, P.R. China
@@
Instituto Avanzado de Cosmología A. C., San Marcos 11 - Atenas 202. Magdalena Contreras, 10720. Ciudad de México, México
Department of Physics and Astronomy, University of Waterloo, 200 University Ave W, Waterloo, ON N2L 3G1, Canada
Waterloo Centre for Astrophysics, University of Waterloo, 200 University Ave W, Waterloo, ON N2L 3G1, Canada
@@
Perimeter Institute for Theoretical Physics, 31 Caroline St. North, Waterloo, ON N2L 2Y5, Canada
Space Sciences Laboratory, University of California, Berkeley, 7 Gauss Way, Berkeley, CA 94720, USA
Max Planck Institute for Extraterrestrial Physics, Gießenbachstraße 1, 85748 Garching, Germany
@@
Department of Physics, Kansas State University, 116 Cardwell Hall, Manhattan, KS 66506, USA
Department of Physics and Astronomy, Sejong University, Seoul, 143-747, Korea
@@
Centre for Astrophysics & Supercomputing, Swinburne University of Technology, P.O. Box 218, Hawthorn, VIC 3122, Australia
CIEMAT, Avenida Complutense 40, E-28040 Madrid, Spain
Department of Physics, University of Michigan, Ann Arbor, MI 48109, USA
@@
Department of Astronomy, Tsinghua University, 30 Shuangqing Road, Haidian District, Beijing, China, 100190

### v3

@@ A.1 Angular bins
Analogously, and are the 3-point and connected 4-point correlation functions of the tracers listed in the superscript evaluated at the separations between and points, respectively. These non-Gaussian higher-point functions are included for completeness but dropped in practice, we reflected this by crossing them out in the expressions.
The expressions for multi-tracer jackknife covariance can be constructed similarly. However, to the best of our knowledge, only the single-tracer jackknife auto-covariances have been used in practice. The shot-noise rescaling values have been tuned on each tracer separately. Therefore the computation of the jackknife model for the other blocks of the full matrix has been unnecessary; the single-tracer expressions are sufficient.
A.2 Projected Legendre
Note that the Legendre projection factors (Eq. 3.2) are the same for all tracers. Accordingly, the multi-tracer covariance is still projected linearly from the angularly binned one:
Therefore the full covariance model is constructed analogously to Eq. A.4:
with the following terms (similar to Eq. 3.4):
The multi-tracer covariance model for Legendre multipoles of the 2PCF is implemented in the code and has been employed for overlapping tracers’ cross-correlation in [39]. We omit the corresponding jackknife expressions as they have not been used.
@@ Appendix B Covariance for the combination of two regions
In this section, we document the procedure to compute the covariance matrix for the combination of two (or more) disjoint regions (volumes) using the covariances in each region. It is fully consistent with the correlation function treatment in the DESI data processing pipeline.
@@ B.1 Angular bins
During the combination of regions labeled by , DESI scripts simply add the total counts of every relevant kind (where each can be or , data or randoms353535I.e. encompasses , , and count arrays if standard BAO reconstruction is not used. After reconstruction we also have – shifted randoms, and the counts relevant to the Landy-Szalay estimator are , , , and .):
The following technical details are important for the strictly correct implementation with pycorr. The counts entering the correlation function estimator (Eq. A.1) need to be normalized properly. Originally the main cause was the possibly different number of galaxies and random points [49, 91]. Now we are often weighting both types of points, and the pair counts weighted by the product of the two individual point weights can be normalized by the product of the sums of the weights363636Exactly for cross-correlations () and for auto-correlations.. pycorr stores both the weighted counts (wcounts) and their normalization (wnorm). It is their ratio, the normalized counts, that enter the correlation function estimator (Eq. A.1). Before adding the pair counts for the two regions, DESI scripts bring counts of different types to the same normalization within each region (via the normalize method373737By default, this method preserves the wnorm of counts.). During the combination of regions wnorm are added as well as wcounts. Therefore the change of normalization within each region alters the relative contributions of the regions to different types of counts, even though it does not affect the region’s correlation function. Therefore this step is important to reproduce. With all count types brought to the same normalization (in each region and therefore in their combination), the wnorm cancel out in the Landy-Szalay estimator (Eq. A.1) and so it is sufficient to substitute only wcounts.
Then the total correlation function (Eq. A.1) is simply the average of the two regions’ correlation functions weighted by counts:
@@
where we neglect the covariance between the correlation functions in different regions, the estimation of which poses extra challenges. We expect this to be a safe approximation for the North and South Galactic caps in the DESI footprint because the separation between galaxies in different caps is significantly larger than the maximum separation for correlation function measurements (200 ).
@@ B.2 Legendre
One can do the reverse approximately with bin-averaged values of the Legendre polynomials383838Or possibly with just the middles of the bins, but this way they end up very much related to already computed projection factors .:
These do not depend on the region. Additional approximation comes from the fact that we limit the multipole index — in this work, we only consider .
With this we can work out the partial derivative of the Legendre moment of the combined correlation function with respect to one in each of the regions in the following steps:
@@
Appendix C Author affiliations
1Center for Astrophysics Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA
@@ Appendix C Author affiliations
2Institute of Physics, Laboratory of Astrophysics, École Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, Chemin Pegasi 51, CH-1290 Versoix, Switzerland
3IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
4Physics Department, Yale University, P.O. Box 208120, New Haven, CT 06511, USA
5Department of Physics & Astronomy, Ohio University, Athens, OH 45701, USA
@@
6Center for Cosmology and AstroParticle Physics, The Ohio State University, 191 West Woodruff Avenue, Columbus, OH 43210, USA
@@
7Department of Astronomy, The Ohio State University, 4055 McPherson Laboratory, 140 W 18th Avenue, Columbus, OH 43210, USA
8Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA
9Physics Dept., Boston University, 590 Commonwealth Avenue, Boston, MA 02215, USA
10University of Michigan, Ann Arbor, MI 48109, USA
@@
11Leinweber Center for Theoretical Physics, University of Michigan, 450 Church Street, Ann Arbor, Michigan 48109-1040, USA
@@
12Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK
@@
13Institute for Computational Cosmology, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK
@@
14Instituto de Física, Universidad Nacional Autónoma de México, Circuito de la Investigación Científica, Ciudad Universitaria, Cd. de México C. P. 04510, México
@@
15Department of Astronomy, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
@@
16Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Menlo Park, CA 94305, USA
17SLAC National Accelerator Laboratory, Menlo Park, CA 94305, USA
18University of California, Berkeley, 110 Sproul Hall #5800 Berkeley, CA 94720, USA
@@
19Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra Barcelona, Spain
20Departamento de Física, Universidad de los Andes, Cra. 1 No. 18A-10, Edificio Ip, CP 111711, Bogotá, Colombia
@@
21Observatorio Astronómico, Universidad de los Andes, Cra. 1 No. 18A-10, Edificio H, CP 111711 Bogotá, Colombia
22Department of Physics, The University of Texas at Dallas, Richardson, TX 75080, USA
@@
23Departament de Física Quàntica i Astrofísica, Universitat de Barcelona, Martí i Franquès 1, E08028 Barcelona, Spain
24Institut d’Estudis Espacials de Catalunya (IEEC), 08034 Barcelona, Spain
@@
25Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (UB), c. Martí i Franquès, 1, 08028 Barcelona, Spain.
@@
26Consejo Nacional de Ciencia y Tecnología, Av. Insurgentes Sur 1582. Colonia Crédito Constructor, Del. Benito Juárez C.P. 03940, México D.F. México
@@
27Departamento de Física, DCI-Campus León, Universidad de Guanajuato, Loma del Bosque 103, León, Guanajuato C. P. 37150, México
28Fermi National Accelerator Laboratory, PO Box 500, Batavia, IL 60510, USA
@@
29Department of Physics, The Ohio State University, 191 West Woodruff Avenue, Columbus, OH 43210, USA
30School of Mathematics and Physics, University of Queensland, 4072, Australia
31NSF NOIRLab, 950 N. Cherry Ave., Tucson, AZ 85719, USA
@@
32Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), FR-75005 Paris, France
@@
33Departament de Física, Serra Húnter, Universitat Autònoma de Barcelona, 08193 Bellaterra (Barcelona), Spain
@@
34Laboratoire de Physique Subatomique et de Cosmologie, 53 Avenue des Martyrs, 38000 Grenoble, France
@@
35Institució Catalana de Recerca i Estudis Avançats, Passeig de Lluís Companys, 23, 08010 Barcelona, Spain
36Department of Physics and Astronomy, University of Sussex, Brighton BN1 9QH, U.K
37Department of Physics & Astronomy, University of Wyoming, 1000 E. University, Dept. 3905, Laramie, WY 82071, USA
@@
38National Astronomical Observatories, Chinese Academy of Sciences, A20 Datun Rd., Chaoyang District, Beijing, 100012, P.R. China
@@
39Instituto Avanzado de Cosmología A. C., San Marcos 11 - Atenas 202. Magdalena Contreras, 10720. Ciudad de México, México
40Department of Physics and Astronomy, University of Waterloo, 200 University Ave W, Waterloo, ON N2L 3G1, Canada
41Waterloo Centre for Astrophysics, University of Waterloo, 200 University Ave W, Waterloo, ON N2L 3G1, Canada
@@
42Perimeter Institute for Theoretical Physics, 31 Caroline St. North, Waterloo, ON N2L 2Y5, Canada
43Space Sciences Laboratory, University of California, Berkeley, 7 Gauss Way, Berkeley, CA 94720, USA
44Max Planck Institute for Extraterrestrial Physics, Gießenbachstraße 1, 85748 Garching, Germany
@@
45Department of Physics, Kansas State University, 116 Cardwell Hall, Manhattan, KS 66506, USA
46Department of Physics and Astronomy, Sejong University, Seoul, 143-747, Korea
@@
47Centre for Astrophysics & Supercomputing, Swinburne University of Technology, P.O. Box 218, Hawthorn, VIC 3122, Australia
48CIEMAT, Avenida Complutense 40, E-28040 Madrid, Spain
49Department of Physics, University of Michigan, Ann Arbor, MI 48109, USA
@@
50Department of Astronomy, Tsinghua University, 30 Shuangqing Road, Haidian District, Beijing, China, 100190